PROBLEM 3 Determine the following in drops per minute using the drop factor provided. 250 ml.ot hours 1. mL/hr packed cls in 2. gtt/min (drop factor: 20 gtt/mL):
Answer #1

Volume= 250 ml

Time= 4 hours= 4*60= 240 minutes.

1.Flowrate (mL/Hr)= Volume(mL)/Time(Hr)= 250/4=62.5 mL/Hr.

2. Flowrate(gtts/min)= Flowrate (mL/Hr)* Drop factor/ Time in minutes= 62.5*20/4*60= 5.2 gtts/min